# supplement.pdf

### v1

@@ Abstract
We observe in Fig. 4B that the distribution is flat for the nearly straight cells of curvature $0.011/R=$ 0.01. Thus, we do not find spatio-orientational coupling for non-curved cells. However, increasing cellular curvature to $0.101/R=$ 0.10 and $0.33$, a peak emerges in the distribution at $\theta_{i}=\gamma_{i}$, see Figs. 4C and D. Consequently, here, the spatio-orientational coupling emerges as described above. This trend is observed even for the smaller division length of $L_{\mathrm{d}}=3$ and $4$ (see Section 2 of the Supporting Information).
To further quantify the degree of spatio-orientational coupling, we fit the Gaussian $a\exp[(\theta-\gamma)^{2}/(2\sigma^{2})]+b$ to each of the unimodal histograms and extract the standard deviation $\sigma$, amplitude $a$, and vertical shift $b$. The dashed lines in Figs. 4C and D confirm the suitability of the fitting function. Figures 4E and F display the resulting standard deviation $\sigma$ and the amplitude $a$, respectively, of the Gaussian fit as a function of the curvature $1/R$. Fitting is performed for $1/R\geq 0.067$. We consider threshold division lengths of $L_{\mathrm{d}}=5$, $4$, and $3$.
For $L_{\mathrm{d}}=5$, we find that the standard deviation $\sigma$ in Fig. 4E increases nearly monotonically with cellular curvature. Large values of $\sigma$ imply less peaked distributions. At the same time, the amplitude of the modulation of the spatio-orientational coupling in Fig. 4F, after an initial increase, drops for elevated curvature. The drop is in line with the decoupling of position and orientation signaled by the standard deviation. Possibly, this is due to the higher isotropy of strongly bent longer cells. In other words, the contour of a long cell at some point closes into a circle, which implies isotropic steric interactions. Consequently, from Figs. 4E and F, we infer that for $L_{\mathrm{d}}=5$ spatio-orientational coupling is largest at intermediate values of curvature, when $\sigma$ is still moderately low and $a$ is pronounced.
Contrarily, for on average shorter cells of division lengths $L_{\mathrm{d}}=4$ and $3$, the standard deviation $\sigma$ in Fig. 4E behaves in a nonmonotonic way with increasing cellular curvature $1/R$. There is no clear trend, in contrast to the behavior for $L_{\mathrm{d}}=5$. Conversely, the amplitude of the Gaussian fit $a$ in Fig. 4F indicates qualitatively the same behavior for all $L_{\mathrm{d}}$. All three lines show a maximum around $0.21/R=$ 0.2. This indicates the most pronounced spatio-orientational coupling at intermediate curvature also for $L_{\mathrm{d}}=4$ and $3$. Our observation suggests that for all considered division lengths, at least at intermediate curvature $1/R$, a spatio-orientational coupling exists. It implies that the normals to the arc-like contours of the microorganisms tend to point radially outward of the colony.
@@
To illustrate this feature and to contrast the organization in colonies of straight and curved cells, we include Fig. 5. There, we plot by dots the center positions $\bm{r}_{i}^{\mathrm{o}}$ of the arcs of all cells, see Fig. 1A, together with the color coding representing cellular orientations. For almost straight microorganisms of curvature $0.011/R=$ 0.01, we hardly observe any such stacking of nearby aligned cells; see Fig. 5A. The spatial organization of the center-of-arc positions $\bm{r}_{i}^{\mathrm{o}}$ is relatively homogeneous and gas-like. However, the situation changes qualitatively with increasing curvature, here to $0.101/R=$ 0.10 and $0.33$ in Figs. 5B and C. With increasing curvature, we find branched filamentous elements of $\bm{r}_{i}^{\mathrm{o}}$. They are a clear signature of the line-like stacking arrangements of neighboring curved cells in the corresponding colonies, in contrast to what is observed for straight cells.
Conclusions
In summary, we address the planar growth of colonies of bent bacteria and other microorganisms. To this end, we consider the persistent curvature of the individual units. These units grow along their arc length. When reaching a certain threshold length, they divide. Only mutual steric interactions that represent volume exclusion between the individuals are included.
@@ Conclusions
Our focus in this initial study is on the fundamental phenomenology underlying and resulting from the expansion of colonies composed of growing and dividing curved microorganisms. We concentrate on the generic effect of volume exclusion. On this basis, extensions into various directions are conceivable. For instance, mechanical regulation of the speed of individual growth [20, 11], diffusive processes [16], or polydispersity of aspect ratios [37] can become important. Along these lines, the role of nutrient supply and its effect on the growth rate represents another interesting aspect and takes the impact of the environment into account [14, 15]. Additional types of mutual interaction can come into play, for example, adhesional forces [38]. Moreover, beyond planar arrangements in monolayered colonies, the transition to bi- and multilayered structures and thus into the third dimension becomes important in reality at some point [39, 40, 41]. Such questions are being addressed for straight cells. Clarifying the role of curvature of the individuals in combination with all these effects implies an interesting subject of research on its own.
Acknowledgements.
We thank the Deutsche Forschungsgemeinschaft (German Research Foundation, DFG) for support through the Heisenberg grant no. ME 3571/4-1. Y.K. acknowledges support from JSPS KAKENHI (grant no. JP23KJ1068). T.K. acknowledges support by the JST FOREST Program (grant no. JPMJFR212T), AMED Moonshot Program (grant no. JP22zf0127009), and JSPS KAKENHI (grant no. JP24H02203).

### v2

@@ Results
We observe in Fig. 4(b) that the distribution is flat for the nearly straight cells of curvature $0.011/R=$ 0.01. Thus, we do not find spatio-orientational coupling for non-curved cells. However, increasing cellular curvature to $0.101/R=$ 0.10 and $0.33$, a broad peak emerges in the distribution around $\theta_{i}=\gamma_{i}$, see Figs. 4(c) and (d). Consequently, here, the spatio-orientational coupling emerges as described above. This trend is observed even for the smaller division length of $L_{\mathrm{d}}=3$ and $4$ (see the supporting material).
To further quantify the degree of spatio-orientational coupling, we fit the circular normal distribution (also known as the von Mises distribution), which is a $2\pi$-periodic analog of the normal distribution [39], to each of the unimodal histograms:
$f(\theta)=\frac{1}{2\pi I_{0}(K)}e^{K\cos(\theta-\gamma)},$ (5)
where $I_{0}(K)$ is the zeroth-order modified Bessel function of the first kind. The dashed lines in Figs. 4(c) and (d) confirm the suitability of the fitting function.
The parameter $K$ characterizes the strength of the modulation and thus the height of the peak of the distribution. Therefore, the larger the value of $K$, the more pronounced the spatio-orientational coupling is. Figure 4(e) displays the resulting dependence of $K$ on the curvature $1/R$ for $1/R\geq 0.067$. The error bars indicating standard deviations in Fig. 4(e) were obtained from the fits to the different realizations of the system, averages of which are displayed as the data points. We again consider threshold division lengths of $L_{\mathrm{d}}=5$, $4$, and $3$.
All three lines feature a maximum around $0.21/R=$ 0.2. Elevated values of $K$ imply more strongly peaked distributions. This suggests that the most pronounced spatio-orientational coupling occurs at intermediate curvature. It implies that the normals to the arc-like contours of the microorganisms tend to point radially outward of the colony. When reducing curvature $1/R$ from the maximum of the curves, $K$ drops, signaling flatter distributions and thus lower spatio-orientational coupling. This trend is in line with observations for straight cells as in Fig. 4(b), where spatio-orientational coupling is absent. Thus, the effect of spatio-orientational coupling is introduced by curvature. As for the decrease of $K$ at large curvatures, possibly, this is due to the higher isotropy of strongly bent cells. In other words, the contour of a strongly curved long cell closes into a circle, which implies more isotropic steric interactions.
@@
To illustrate this feature and to contrast the organization in colonies of straight and curved cells, we include Fig. 5. There, we plot by dots the center positions $\bm{r}_{i}^{\mathrm{o}}$ of the arcs of all cells, see Fig. 1(a) and the inset of Fig. 5(a), together with the color coding representing cellular orientations. For almost straight microorganisms of curvature $0.011/R=$ 0.01, we hardly observe any such stacking of nearby aligned cells; see Fig. 5(a). The spatial organization of the center-of-arc positions $\bm{r}_{i}^{\mathrm{o}}$ is relatively homogeneous and gas-like. However, the situation changes qualitatively with increasing curvature, here to $0.101/R=$ 0.10 and $0.33$ in Figs. 5(b) and (c). With increasing curvature, we find branched filamentous elements of $\bm{r}_{i}^{\mathrm{o}}$. They are a clear signature of the line-like stacking arrangements of neighboring curved cells in the corresponding colonies, in contrast to what is observed for straight cells. Curved cells can be stacked like cups into each other to save space, whereas this is not necessary for straight objects that can form orientationally ordered domains for space-efficient arrangement.
Conclusion
In summary, we address the planar growth of colonies of bent bacteria and other microorganisms. To this end, we consider persistent curvature of the individual units. These units grow along their arc length. When reaching a certain threshold length, they divide. Only mutual steric interactions that represent volume exclusion between the individuals are included.
@@ Conclusion
Our focus in this initial study is on the fundamental phenomenology underlying and resulting from the expansion of colonies composed of growing and dividing curved microorganisms. We concentrate on the generic effect of volume exclusion. On this basis, extensions into various directions are conceivable. For instance, mechanical regulation of the speed of individual growth [20, 11], diffusive processes [16], polydispersity of aspect ratios [40], or anisotropic friction [16, 36, 37] can become important. Additional types of mutual interaction can come into play, for example, adhesional forces [41]. Moreover, beyond planar arrangements in monolayered colonies, the transition to bi- and multilayered structures and thus into the third dimension becomes important in reality at some point [42, 43, 44].
Besides, one may think not only about extensions that concern the influence of properties related to the cells themselves. For instance, the role of nutrient supply and its effect on the growth rate represents another interesting aspect and takes the impact of the environment into account [14, 15]. Along these lines, the mechanical properties of the substrate can play a notable role as well. Soft substrates deform under the action of the stress induced by living biological cells [45, 46]. Resulting interactions via deformations of elastic films are long-ranged. They can affect or induce mutual orientational order [47, 48].
Many of the just-listed aspects are being addressed for straight cells. Clarifying the role of curvature of the individuals in combination with all these effects implies an interesting subject of research on its own. For instance, the shape of the individual curved objects becomes increasingly isotropic during growth towards a straight circle. In this way, during the growth cycle of each individual, anisotropy in friction and in mutual long-ranged interaction with and through the substrate may vary.
Author Contributions
A.M.M. designed the research; Y.K. performed the simulations and analyzed the numerical data; T.K. and A.M.M. assisted with the numerical implementation and data analysis; all authors discussed and interpreted the results; all authors wrote the paper.
Acknowledgments
We thank the Deutsche Forschungsgemeinschaft (German Research Foundation, DFG) for support through the Heisenberg grant no. ME 3571/4-1. A.M.M. further thanks the DFG for support through the research grant ME 3571/12-1. Y.K. acknowledges support from JSPS KAKENHI (grant no. JP23KJ1068). T.K. acknowledges support by the JST FOREST Program (grant no. JPMJFR212T), AMED Moonshot Program (grant no. JP22zf0127009), and JSPS KAKENHI (grant no. JP24H02203).
Supporting Material
Supporting material is attached below.